Three forces, F_1 = 40 N, F_2 = 20 N and F_3 = 10
Answer #1

normal force = mg - F3 - F2*sin30 = 4*9.8 - 10 - 20/2 = 19.2 N

a)

friction force = uk*N = 0.12*19.2 = 2.304 N

b)

horizontal force on block

F = F1 + F2*cos30 - friction force = 40 + 20*0.8660254 - 2.304 = 55.016 N

acceleration of the block in x axis = a = F/m = 55.016/4 = 13.75 m/s^2